OpenStudy (anonymous):

15x - 2 = 4x - 13 A. x = -1 B. x = 15/19 C. x = 1 D. x = -11/19 2. Simplify: 2/3x = 54 A. x = 36 B. x = 18 C. x = 81 D. None of the above @Fransisco_w @sangya21

OpenStudy (anonymous):

15x - 2 = 4x - 13 15x -4x = -13 + 2 11x = - 11 x = -1

OpenStudy (anonymous):

For the first question you want to get x on one side then isolate it. do that by subtracting 4x from both sides and adding 2 to both sides as well. you get this after that 11x = 11 now divide the 11 out to get x =1

OpenStudy (anonymous):

1 OR -1 @Fransisco_w

OpenStudy (anonymous):

2/3x = 54 2/54 = 3x 1/27 = 3x x = 3 * 1/27 x = 1/9

OpenStudy (anonymous):

Sangya you way wrong on that answer

OpenStudy (anonymous):

To get rid of of 2/3 you multiply it by 3/2 on both sides which would look like this... x = 54 x 3/2 x = 162/2 x = 81

OpenStudy (anonymous):

oh and its -1 for the first question :p